Calculate the number of formula units of Li2CO3 present in a 0.500g sample of lithium carbonate. Li=7, C=12, O=16


1
Expert's answer
2022-02-09T04:54:04-0500

Solution:

Calculate the molar mass of Li2CO3:

Molar mass of Li2CO3 = 2×Ar(Li) + Ar(C) + 3×Ar(O) = 2×7 + 12 + 3×16 = 74 (g/mol)


Calculate the moles of Li2CO3:

Moles of Li2CO3 = (0.500 g Li2CO3) × (1 mol Li2CO3 / 74 g Li2CO3) = 0.0067568 mol Li2CO3


One mole of any substance contains 6.022×1023 formula units.

Hence,

Number of formula units of Li2CO3 = (0.0067568 mol) × (6.022×1023 formula units / 1 mol) = 4.069×1021 formula units Li2CO3


Answer: The number of formula units of Li2CO3 is 4.069×1021